Grade 12th passInorganic Chemistry

Bond angle order of no2, no2-, no2+,no3- in decreasing order

Zafar Imam Ansari
7 Years agoGrade 12th pass

3 Answers

Arun
7 Years ago
NO2- 3- 2 2+ 

For NO2-, the molecule is bent but we have 1 lone pair on nitrogen. Lone pair electron repulsion is more than the bonding electrons so the bond angle will be less than 120o.
In NO3-, there is no lone pair of electrons, all are bonding pairs leading to an ideal bond angle of 120o.
In NO2the one lone electron exerts a less repulsion than a lone pair of electrons, so two bonding oxygen atoms are able to spread out more leading to bond angle greater than the ideal of 120o.
In NO2+, there is no lone pair, only bond pairs so the molecule is linear leading to a bond angle of 180o.
Tony
7 Years ago
An easy way to calculate bond order is:
 
Bond Order=Valancy of peripheral atom + (Charge on whole atom) / No. Of peripheral atoms. 
 
Example :
For No2-, peripheral atom here is oxygen
Bond order = 2 + (-1)/2 
                    =2 - 0.5 = 1.5
 
Kushagra Madhukar
6 Years ago
Hello student
 
The pictorial representation of the species mentioned and their respective angle in ascending order is as shown.
Let’s analyse each specie individually
NO2- → In this the hybridisation of central N atom is sp2, which accounts for 120o bond angle, but since the repulsion given by the lone pair of electrons is greater than the bond pair electrons, the O-N-O bond angle decreases and attains a value lesser than 120o
NO3- → In this the hybridisation of central N atom is sp2 and also there is no lone pair and only bond pair of electrons(which also undergo resonance and equalises bond order), therefore the angle between O-N-O remains unchanged and is equal to 120o
NO2 → In NO2, instead of having a lone pair of electrons it only has a single elecron in its lone pair, which gives lesser repulsion to the O-N-O bond, even lesser than the bond pair – bond pair repulsion, therefore the O-N-O bond readjusts and its angle increases and attains a value greater than 120o
NO2+ → In NO2+, the central N atom is sp hybridised and there is no lone pair on N atom, thus the angle between O-N-O bond becomes 180o
 
Therefore the order of angle is- NO2- 3- 2 2+
 
Regards,
Kushagra